Leave answers as fractions. Indicate whether it is an add or
multiply.
You are dealt a hand of five cards from a standard deck of 52 cards.
Find the probability of selecting a three or a club?

Found 2 solutions by greenestamps, ikleyn:
Answer by greenestamps(13195)



"OR" means ADD the probabilities.

In this example you need to be careful, because "three" and "club" overlap.

P(three OR club) = P((three) OR (club that is not a three)) = 4/52 + 12/52 = 16/52 = 4/13.

Of course, you could calculate the probability as

P(three OR club) = P((club) OR (three that is not a club)) = 13/52 + 3/52 = 16/52 = 4/13.

Finally, a technique that is often required for more complicated problems is to calculate the probability that it is a club or that it is a three and then subtract the probability that it is both -- because you have counted that probability twice.

P(three OR club) = P(three) + P(club) - P(three of clubs) = 4/52 + 13/52 = 1/52 = 16/52 = 4/13


Answer by ikleyn(52754)

.

As the problem is worded,  printed,  submitted,  formulated and presented in the post,  its meaning is dark and unclear.